OpenStudy (anonymous):

A baseball with a mass of 0.145 kilograms collides with a bat at a velocity of 44 meters/second. The ball bounces off the bat with a speed of 49 meters/second in the opposite direction. What is the impulse

OpenStudy (souvik):

hint:impulse = change in momentum...

OpenStudy (anonymous):

A. 0.72 newton seconds B. 4.32 newton seconds C. 7.25 newton seconds D. 8.23 newton seconds E. 9.42 newton seconds

OpenStudy (souvik):

change in velocity is 49-(-44)=93 m/s change in momentum = 0.145*93=13.485 Ns but it is not in options.